For the random variable X with density function
f(x)=[2e^-2x ; x≥0
[0 ; <0
Find
i) the mean μand variance σ²;
ii) P[|x–μ|≥1] .
Use Chebyshev’s inequality to obtain an upper bound on P[ |X − μ| ≥1]and compare
with the result obtained in part (ii).
1
Expert's answer
2019-04-02T11:34:01-0400

"\mu=\int_{0}^{\infty}2xe^{-2x}dx=1/2\\
\sigma^{2}=\int_0^{\infty}2(x-1/2)^{2}e^{-2x}dx=1/4\\
P(|X-\mu|>=1)=1-P(|X-\mu|<1)=1-\int_{0}^{3/2}2e^{-2x}dx=e^{-3}=0.0498\\"

By Chebyshev's inequality: "P(|X-\mu|>=1)=P(|X-\mu|>=2\sigma)<=1/4=0.25."